OpenStudy (jojo12g):

What value of x makes the equation true? –4.9 = –5.7 + x A. x = –10.6 B. x = –0.8 C. x = 0.8 D. x = 10.6

OpenStudy (alexandervonhumboldt2):

-4.9+5.7=x x=5.7-4.9=?

OpenStudy (alexandervonhumboldt2):

explain your answer @amber12362

OpenStudy (derpisthewordorisit):

^

OpenStudy (anonymous):

Oh ok

OpenStudy (derpisthewordorisit):

its A

OpenStudy (derpisthewordorisit):

It's -10.6, you forgot the negative part @AlexandervonHumboldt2

OpenStudy (ddos_dragon):

–5.7 +10.6 = 4.9

OpenStudy (alexandervonhumboldt2):

what part? are you laughing on me?

OpenStudy (derpisthewordorisit):

I'm not laughing at you, you forgot next to 5.7

OpenStudy (alexandervonhumboldt2):

–4.9 = –5.7 + x x=-4.9+5.7 x=5.7-4.9=0.8

OpenStudy (anonymous):

 It's actually C babe.